factorise x^2+px+qx+pq
factorise m^2-4
write as a single fraction in its simplest form 2/x-4-1/x+3

Answer by ewatrrr(24785)

x^2+px+qx+pq = x(x+p) + q(x+p) = (x+q)(x+p)
m^2-4 = (m-2)(m+2)